Need help with this word problem please:
A washer and a dryer cost $841 combined. The washer costs $91 more than the dryer. What is the cost of the dryer?
Here is my try:
w+d=841
w+91+d=841?

Found 2 solutions by Earlsdon, jsmallt9:
Answer by Earlsdon(6294)

Nice try but not quite correct!
1) W+D = 841 This part is ok.
2) W = D+91 "The washer costs $91 more than the dryer"
Substitute the W of equation 2) into equation 1) and solve for D.
(D+91)+D = 841 Simplify.
2D+91 = 841 Subtract 91 from both sides.
2D = 750 Finally, divide both sides by 2.
D = $375

Answer by jsmallt9(3758)

Here is my try:
w+d=841
w+91+d=841

The first equation is good. The second equation ... not so much.

The second equation should state the relationship between the cost of the dryer and washer. Since the "washer costs $91 more than the dryer" we can use:
w=91+d

So our two equations are:
w+d=841
w=91+d
Using the second equation, since it is already solved for w, we will substitute for w into the first equation:
(91+d)+d=841
Simplifying we get:
91 + 2d = 841
Subtracting 91 from both sides:
2d = 750
Dividing both sides by 2:
d = 375
So the dryer costs $375. To find the cost of the washer we can substitute this number for d in one of the 2 equations. Using the second equation we get:
w = (375) + 91
Adding we get:
w = 466
So the washer costs $466